the sqroot of (3x+10)^2= (5-2x)^2
Answer by edjones(8007)

3x+10=(5-2x)^2 remove the square sign from the left side.
...=4x^2-20x+25
4x^2-23x+15=0
4*15=60; The 2 factors of 60 whose sum is -23 are -3 and -20.
4x^2-3x-20x+15=0
x(4x-3)-5(4x-3)=0 factoring by grouping.
(x-5)(4x-3)=0
x=5, x=3/4
.
Ed